OpenStudy (kinggeorge):

@nincompooop Let \(\vec{v}=(v_1,v_2,v_3)\) and \(\vec{w}=(w_1,w_2,w_3)\). We want to find \(\vec{v}\times\vec{w}\), show it's perpendicular, and show that \(\vec{v}\times\vec{w}=-\vec{w}\times\vec{v}\).

OpenStudy (kinggeorge):

In matrix form, the cross product is\[\left|\begin{matrix} \vec{i}&\vec{j}&\vec{k}\\ v_1&v_2&v_3\\ w_1&w_2&w_3\\ \end{matrix}\right|=(v_2w_3-v_3w_2)\vec{i}-(v_1w_3-w_1v_3)\vec{j}+(v_1w_2-w_1v_2)\vec{k}\]Taking dot products with \(\vec{v}\) and \(\vec{w}\), we get that the cross product is perpendicular.

OpenStudy (kinggeorge):

As for the second thing we want to prove, recall that when taking the determinant, elementary row operations can change it in a subtle way. In particular, swapping two rows, multiplies the determinant by \(-1\). So if we swap the bottom two rows, we get that\[\left|\begin{matrix} \vec{i}&\vec{j}&\vec{k}\\ v_1&v_2&v_3\\ w_1&w_2&w_3\\ \end{matrix}\right|=-\left|\begin{matrix} \vec{i}&\vec{j}&\vec{k}\\ w_1&w_2&w_3\\ v_1&v_2&v_3\\ \end{matrix}\right|=-\vec{w}\times\vec{v}\]

OpenStudy (kinggeorge):

I'm actually TA'ing a linear algebra class this semester, so if you feel like shooting me more related questions I'll do my best to answer them.

OpenStudy (nincompoop):

yeah, I will. thank you! I need some math fundamentals for my electromagnetism in physics class

OpenStudy (kinggeorge):

Well, it's bedtime for me. I'm making an effort to be a little more present on this website though, and if you message me, it should get forwarded to my email. So if I'm not on when you have a question, and you for some reason desperately want me to answer it, you can try to message me.